OpenStudy (faith):

how to find the antiderivative of xroot ofx^2 +1

OpenStudy (anonymous):

Wait, is this \[\sqrt[x]{x^2+1}\] or \[x * \sqrt{x^2+1}?\]

OpenStudy (faith):

x.sqrt(X^2 +1)

OpenStudy (anonymous):

If it's the second option (which I think it is), then you have ∫ x √(x^2+1) dx. Let u = x^2+1, and you end up with du = 2x dx. Hence, dx = du/2x, and when you plug it into your integrand, you end up with ∫ 1/2 * √u du. You should be able to solve it after that. :)